A brave child decides to grab on to an already spinning merry-go-round. The child is initially at rest and has a mass of 26.5 kg. The child grabs and clings to a bar that is 1.45 m from the center of the merry-go-round, causing the angular velocity of the merry-go-round to abruptly drop from 55.0 rpm to 19.0 rpm. What is the moment of inertia of the merry-go-round with respect to its central axis?
Answer #1
  • from angular momentum conservation;
    I *55= (I+26.5* 1.45^2) 19
    => 2.895I = I + 55.716
    => I = 29.40 kg-m2
Answer #3

石om the Cons ervaho幻6f ang uloy momentum